OpenStudy (anonymous):

if h(x) - (f o g) (x) and h(x) = |9x +10|, find functions f and g

OpenStudy (freckles):

Example: If we are given h(x)=(2x)^3 we could say the inner function g(x)=2x and the outer function f(x)=x^3

OpenStudy (freckles):

verifying; h=f(g(x))=f(2x)=(2x)^3

OpenStudy (freckles):

So if have h(x)=|9x+10| can you identify the outer function?

OpenStudy (anonymous):

Would the outer function be 1x in this instance?

OpenStudy (freckles):

what about the | |

OpenStudy (freckles):

\[h(x)=|inside|\]

OpenStudy (freckles):

so the the outside has to be | |

OpenStudy (freckles):

so the outer function f(x)=|x| The only thing left is to identify the inner function g(x)

OpenStudy (anonymous):

With g(x) being 9x+10

OpenStudy (freckles):

yep

OpenStudy (anonymous):

Absolute values continue to trip me up. Thank you!

OpenStudy (freckles):

np

OpenStudy (freckles):

and absolute value is just another function